# Towards Next-Generation Urban Decision Support Systems through AI-Powered Generation of Scientific Ontology using Large Language Models - A Case in Optimizing Intermodal Freight Transportation

### v1

@@ Abstract
The incorporation of Artificial Intelligence (AI) models into various optimization systems is on the rise. Yet, addressing complex urban and environmental management problems normally requires in-depth domain science and informatics expertise. This expertise is essential for deriving data and simulation-driven for informed decision support. In this context, we investigate the potential of leveraging the pre-trained Large Language Models (LLMs). By adopting ChatGPT API as the reasoning core, we outline an integrated workflow that encompasses natural language processing, methontology-based prompt tuning, and transformers. This workflow automates the creation of scenario-based ontology using existing research articles and technical manuals of urban datasets and simulations. The outcomes of our methodology are knowledge graphs in widely adopted ontology languages (e.g., OWL, RDF, SPARQL). These facilitate the development of urban decision support systems by enhancing the data and metadata modeling, the integration of complex datasets, the coupling of multi-domain simulation models, and the formulation of decision-making metrics and workflow. The feasibility of our methodology is evaluated through a comparative analysis that juxtaposes our AI-generated ontology with the well-known Pizza Ontology employed in tutorials for popular ontology software (e.g., protégé). We close with a real-world case study of optimizing the complex urban system of multi-modal freight transportation by generating anthologies of various domain data and simulations to support informed decision-making.
@@ 1 Introduction
The global urban population has surged from 1.01 billion in 1960 to 4.52 billion in 2022 and is expected to reach 6.9 billion, or approximately 70% of the global population, by 2050 (Hoornweg and Pope, 2017). The accelerated pace of global urbanization underscores the urgent need for effective management and optimization of urban systems. The goals are to promote urban livability through efficient use of resources (Habitat, 2020) and reduce carbon footprint worldwide for more sustainable economic growth under the mounting pressure of climate change (Moran et al., 2018).
@@
Over recent decades, academia, industry, and governmental agencies have significantly invested in developing smart city applications (Angelidou et al., 2018). These efforts aim to foster smart and sustainable cities to enhance urban planning and management and optimize city operations, through the integration of cutting-edge computing and communication technologies (Bibri and Krogstie, 2017). The synergy of these emerging technologies, specifically the Internet of Things (IoT), digital twins, and artificial intelligence (AI), alongside advanced simulations across various urban domains (including transportation, water supply, buildings, and energy systems), fosters a more informed, reliable, and time-sensitive approach to complex urban decision-making (Rane, 2023).
Despite the successes of past smart city applications, challenges persist in optimizing complex urban systems, which involve intricate interactions among various subsystems (McPhearson et al., 2016). A prime example is the intermodal freight transportation system, spanning vast areas and governed by dynamic, multi-scale processes with broad socio-economic and environmental impacts (Heinold and Meisel, 2020).
At the methodological level, the development of effective smart city applications faces formidable challenges, particularly due to the need for interdisciplinary knowledge to create scenario-based ontologies that represent these dynamics accurately (Matei et al., 2021). These ontologies are vital for improving problem understanding, formulating objectives, and designing optimization workflows. From a technical standpoint, integrating diverse urban data and coupling models across domains is crucial for generating both model and data-driven insights, essential for informed decision-making in urban digital twins and decision support systems (Chen et al., 2018). This synthesis supports informed decision-making, making it indispensable in the creation of urban digital twins and decision support systems. Yet, it is often constrained by the laborious nature of the work and the extensive interdisciplinary knowledge required to create and integrate scenario-based ontologies for addressing specific urban challenges.
@@
Recent progress in generative models, notably large language models (LLMs) like GPT-3 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), and Google Gemini (Team et al., 2023) has significantly propelled the research and development of autonomous agents and systems forward. In light of these developments, this paper explores the potential of LLMs to leverage their advanced understanding, reasoning, and generative abilities in human language. Our goal is to embed domain knowledge into scientific ontologies for informed decision support in boosting smart city functionalities. By employing the ChatGPT API as a core reasoning tool, we outline a detailed workflow that incorporates (a) scholarly databases, (b) natural language processing tools, (c) structured prompts from the Methontology workflow, and (d) pre-trained LLM.
We aim to automate the creation of scenario-based ontologies through the analysis of research articles and technical documents on urban data and simulations. Our approach yields knowledge graphs and data dictionaries in standard ontology languages like OWL, RDF, and SPARQL. It enhances aspects of urban decision support systems including data modeling, dataset integration, simulation model linking, and the formulation of decision-making criteria. Further, we evaluate our method’s effectiveness by comparing it with the Pizza Ontology, used in ontology software tutorials such as Protégé. The paper culminates in a case study on optimizing intermodal freight transportation in urban environments, showcasing how our ontologies enable informed decision-making.
@@ 2 Literature Review
Ontology and knowledge graphs are foundational elements in the realm of semantic web and knowledge management, offering structured ways to represent knowledge about the world. Defined as “a formal, explicit specification of a shared conceptualisation” by Noy and Hafner (1997), ontology, in a computational sense, is a formal representation of knowledge as a set of concepts within a domain, and the relationships between those concepts (Gruber, 1993). It provides a shared vocabulary for researchers to annotate, share, and reuse their knowledge across different systems and domains (Chen et al., 2018). Knowledge graphs further extend this concept by creating large networks of entities and their interrelations, encapsulated in graph structures, which enable complex queries and analytics, making knowledge accessible and actionable (Ji et al., 2021).
2.1 Ontology for Smart Cities
@@
In smart city applications, ontology and knowledge graphs play pivotal roles in enhancing urban management, decision support, and the optimization of urban systems. They facilitate the integration of diverse data sources, provide a basis for interoperability between different systems and platforms, and support the reasoning processes needed for intelligent decision-making (Chen et al., 2018). By mapping out the complex relationships between various urban elements such as transportation networks, utilities, and services, these tools help in synthesizing data into actionable insights, thereby improving efficiency, sustainability, and the quality of urban life (Bibri and Krogstie, 2017).
@@
Furthermore, ontology and knowledge graphs are crucial in the development of web-based decision support systems (Spoladore and Pessot, 2021), digital twins (Petrova-Antonova and Ilieva, 2021), databases (Balasubramani et al., 2016), AI applications (Bergman et al., 2018), and software dedicated to smart city management (Shi et al., 2023). They provide the underlying architecture for these technologies to dynamically model and simulate the real-world urban environment, enabling cities to anticipate problems, optimize operations, and engage more effectively with citizens (Kuster et al., 2020). Through the use of these semantic technologies, smart city applications, such as digital twin and online decision support systems, can evolve into more adaptive, responsive, and intelligent software applications, harnessing the full potential of the online big urban data to meet the challenges of complex urban system management and optimization (Weil et al., 2023).
2.2 Ontology Design and Development
Methodologies for constructing scientific ontologies employ both detailed and systemic approaches to address the multifaceted nature of ontology development (Cristani and Cuel, 2005). At the micro level, methodologies emphasize the granular aspects of ontology creation, such as the formalization of axioms and design decisions influenced by the chosen representation language (Guizzardi, 2005). Conversely, macro-level approaches provide a comprehensive systems perspective, focusing on the development process through information systems and IT lens (Merali, 2006).
Key methodologies in the field include older frameworks like Methontology and On-To-Knowledge, alongside more contemporary ones that focus on the re-usability and flexibility of ontology, such as NeON.
@@ 2.2 Ontology Design and Development
    The On-To-Knowledge methodology (OTKM) represents a structured approach to ontology design, aimed at enhancing knowledge management within organizations. It focuses on converting implicit company knowledge into an explicit, structured form, facilitating better decision-making and process optimization (Sure et al., 2004). This method emphasizes early stakeholder involvement, iterative development, and the use of formal ontologies to ensure that the captured knowledge is both accurate and usable in automated systems, making it a cornerstone for developing semantic applications in various domains.
@@
    The NeON methodology caters to the evolving demands of ontology engineering, supporting the reuse, reengineering, and integration of ontological resources (Gómez-Pérez and Suárez-Figueroa, 2009). By offering a diverse set of strategies and tools, NeON facilitates the development of robust ontologies and ontology networks, addressing complex semantic frameworks (Suárez-Figueroa et al., 2015). This methodology is particularly beneficial for projects requiring collaborative and distributed development, making it a pivotal resource for advancing semantic technologies and their applications in various domains.
@@
2.3 Traditional Method for Ontology Implementation
At the implementation level, a variety of software tools for ontology creation, visualization, manipulation, and editing are web-based or desktop-based and are widely utilized by knowledge engineers and domain scientists alike. Among these, Protégé (Noy et al., 2001), OntoSoft (Gil et al., 2016), Hozo (Mizoguchi et al., 2007), and OntoGraf (Falconer, 2010), yEd Graph Editor (Sedlmeier and Gogolla, 2016), Visual Understanding Environment (VUE) (Saigal and Kumar, 2005), and CmapTools (Cañas et al., 2004) have emerged as notable options in recent years. When choosing these tools, ontology developers must consider several critical factors, including the capability for reusing existing ontologies, comprehensive documentation, as well as the ability to export and import data across diverse formats, views, and libraries (Dudáš et al., 2018). These considerations are essential for ensuring the effective application and integration of ontological tools in scientific and engineering workflows.
Implementing ontology with specialized management software is a time-consuming and complex process, largely due to the detailed work required to represent domain-specific knowledge accurately and the expertise needed to effectively utilize these tools (Botzenhardt et al., 2011). The development of ontology is inherently iterative, involving repeated cycles of testing, evaluation, and refinement. Each cycle can reveal new requirements or inconsistencies, necessitating a return to previous stages, which extends the duration of the project (Fernández-López and Gómez-Pérez, 2002). Given these challenges, there is a pressing need to harness emerging AI technologies to automate aspects of ontology implementation, potentially streamlining the process and reducing the manual effort involved.
@@ 2.4 Semantic Web Standards
After being implemented, scientific ontologies are typically stored and documented using common semantic web languages that adhere to the semantic web standards, including RDF (Resource Description Framework), OWL (Web Ontology Language), and SPARQL (SPARQL Protocol and RDF Query Language) (Allemang and Hendler, 2011). These standards are crucial for structuring, storing, and querying data effectively. They enable the structured representation and interconnected querying of data across diverse systems, facilitating seamless data integration and accessibility on the Semantic Web.
@@
Together, these semantic web languages form the backbone of semantic web applications, allowing for robust and sophisticated handling of scientific resources, such as the sharing of scientific datasets and modes, across the internet. Many smart city applications depend on semantic web technologies to discover and integrate essential data, thereby facilitating informed urban decision-making and optimizations. Consequently, these technologies play a pivotal role in the development of numerous urban decision support systems. Therefore, it is imperative that future AI-powered techniques for generating ontologies also adhere to established Semantic Web Standards. This alignment will ensure compatibility and enhance the effectiveness of data integration across various platforms and systems.
@@ 2.5 AI-powered autonomous system
The swift evolution of AI has catalyzed the emergence of autonomous agents (e.g., industrial robots and autonomous vehicles), also referred to as intelligent agents (Brustoloni, 1991), which are sophisticated computer systems designed to undertake tasks and make decisions autonomously, with little to no human oversight (Wooldridge and Jennings, 1995). These agents have demonstrated significant potential across diverse fields such as robotics, healthcare, transportation, and finance, showcasing benefits like enhanced efficiency, diminished errors, and the capability to manage intricate tasks that might be impractical or exceedingly laborious for humans (Stone et al., 2022). Diverging from automated systems, such as automated irrigation systems and home automation systems, that execute predefined tasks based on specific sets of commands, inputs, rules, and environmental conditions, autonomous agents exhibit higher complexity, better adaptability to their surroundings, and the ability to make informed decisions and content generation from collected data (Albrecht and Stone, 2018).
@@
The capability of most existing rule- or algorithm-based reasoning cores is confined to certain or closed environments, limiting their reasoning capacity for tasks that involve creating new content or computer codes. With the advancement of generative AI models, LLMs like GPT-3 (Brown et al., 2020) and GPT-4 (Sun et al., 2023), has significantly propelled forward the research and development of autonomous agents. These models exhibit a profound comprehension of human natural language, enabling them to adeptly perform tasks across various domains including reasoning, creative writing, code generation, translation, and information retrieval (Li and Ning, 2023).
With recent advancements in prompt engineering and tuning techniques, LLMs are increasingly being integrated as the reasoning core to enable AI-powered assistants or autonomous agents capable of performing specialized tasks, such as generating content within specific domain contexts, automating code generation, and synthesizing data (Shin et al., 2023). This integration highlights the transformative impact of LLMs in enhancing automated processes and driving innovation across various fields, shedding light on an autonomous approach for generating scientific ontology.
2.6 Challenge and Knowledge Gap

### v3

# Towards Next-Generation Urban Decision Support Systems through AI-Powered Construction of Scientific Ontology using Large Language Models - A Case in Optimizing Intermodal Freight Transportation

@@ Abstract
The incorporation of Artificial Intelligence (AI) models into various optimization systems is on the rise. However, addressing complex urban and environmental management challenges often demands deep expertise in domain science and informatics. This expertise is essential for deriving data and simulation-driven insights that support informed decision-making. In this context, we investigate the potential of leveraging the pre-trained Large Language Models (LLMs) to create knowledge representations for supporting operations research. By adopting ChatGPT-4 API as the reasoning core, we outline an applied workflow that encompasses natural language processing, methontology-based prompt tuning, and Generative Pre-trained Transformer (GPT), to automate the construction of scenario-based ontologies using existing research articles and technical manuals of urban datasets and simulations. From these ontologies, knowledge graphs can be derived using widely adopted formats and protocols, guiding various tasks towards data-informed decision support. The performance of our methodology is evaluated through a comparative analysis that contrasts our AI-generated ontology with the widely recognized Pizza Ontology, commonly used in tutorials for popular ontology software. We conclude with a real-world case study on optimizing the complex system of multi-modal freight transportation. Our approach advances urban decision support systems by enhancing data and metadata modeling, improving data integration and simulation coupling, and guiding the development of decision support strategies and essential software components.
@@ 1 Introduction
The global urban population has surged from 1.01 billion in 1960 to 4.52 billion in 2022 and is expected to reach 6.9 billion, or approximately 70% of the global population, by 2050 (Hoornweg and Pope, 2017). The accelerated pace of global urbanization underscores the urgent need for effective management and optimization of urban systems. The goals are to promote urban livability through efficient resource allocation (Habitat, 2020) and reduce the global carbon footprint, thereby fostering sustainable economic growth amid increasing climate change pressures (Moran et al., 2018). Over recent decades, academia, industry, and governmental agencies have significantly invested in developing smart city applications (Angelidou et al., 2018). These efforts aim to foster smart and sustainable cities to enhance urban planning and management, optimizing city operations through the integration of cutting-edge computing and communication technologies (Bibri and Krogstie, 2017). The synergy of these emerging technologies, specifically the Internet of Things (IoT), digital twins, and artificial intelligence (AI), alongside advanced simulations across various urban domains (including transportation, water distribution, buildings, and energy systems), promotes a more informed, reliable, and time-sensitive approach to complex urban decision-making (Rane, 2023; Xu et al., 2023b). The approach leverages large volumes of urban data and simulation outputs, combined with expert knowledge and experiences, to produce holistic and resilient strategies that enhance the efficiency of urban management and city operations (Xu et al., 2023a; Weil et al., 2023a).
Despite the successes of past smart city applications, challenges persist in optimizing complex urban systems, which involve intricate interactions among various subsystems (McPhearson et al., 2016). A prime example is the intermodal freight transportation system, spanning vast areas and governed by dynamic, multi-scale processes with broad socioeconomic and environmental impacts (Heinold and Meisel, 2020). Developing holistic solutions to optimize such a system often requires in-depth domain expertise and laborious efforts to conceptualize the problem, formulate decision support strategies, identify relevant data sources and variables that represent essential processes and dynamics, and create a comprehensive data provisioning system (e.g., database and repository) through the integration of diverse urban data and coupled simulation outputs across multiple domains. This data provisioning system is crucial for providing information to analytical and optimization models, as well as decision support workflows, to generate both model and data-driven insights essential for informed decision-making in urban digital twins and decision support systems.
In this regard, many previous studies have taken a knowledge engineering approach, focusing on constructing scenario-based ontologies to conceptualize complex processes in decision support problems (Matei et al., 2021). They create knowledge graphs from these ontologies to support data integration and simulation coupling, thereby guiding the development of efficient data provisioning systems (Chen et al., 2018). This ontology-driven approach is vital for improving problem understanding, identifying objectives, and designing optimization workflows (Wang et al., 2019; Kornyshova and Deneckère, 2010). However, this approach is often constrained by the domain-specialized nature of the work and the extensive interdisciplinary knowledge required to construct scenario-based ontologies for addressing specific urban challenges. In this regard, there is a need to develop an intelligent approach that leverages cutting-edge AI techniques to automatically collect, compile, and interpret interdisciplinary domain knowledge from text documents. This approach should generate scenario-based knowledge representations, such as ontologies and knowledge graphs, to facilitate the development of decision support strategies and data provisioning systems for optimizing complex urban systems.
@@
Recent progress in generative models, notably large language models (LLMs) like GPT-3 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), and Google Gemini (Team et al., 2023), has significantly propelled the research and development of AI agents and autonomous systems forward. In light of these developments, this paper explores the potential of LLMs to leverage their advanced understanding, reasoning, and generative abilities in human language to generate knowledge representations for operations research. By employing the ChatGPT-4 API as a core reasoning tool, we outline a detailed workflow that incorporates (a) scholarly databases, (b) natural language processing tools, (c) structured prompts from the Methontology workflow, and (d) pre-trained LLM. The workflow is designed to automate the construction of scenario-based ontologies through the analysis of research articles and technical documents on recent emerging urban data and simulations. The AI-generated ontologies expand the domain ontology by adding new data sources, variables, and simulation capabilities. From these ontologies, we can drive knowledge graphs in standard formats, such as RDF, using knowledge engineering software and tool. The derived knowledge graphs enhance various aspects of urban decision support systems by enhancing data and metadata modeling, dataset integration, simulation coupling, and the formulation of decision support strategies. Further, we evaluate our method’s effectiveness by comparing its ontology output with the Pizza Ontology, which is often used as the sample data in tutorials of ontology software, such as Protégé. The paper culminates in a case study on optimizing intermodal freight transportation in urban environments, showcasing how our AI-generated ontologies and knowledge graphs are used to guide the integration, management, retrieval, and interpretation of complex multi-dimension datasets, as well as formulation of decision support strategies to enable informed decision optimization.
@@ 2 Literature Review
Ontologies and knowledge graphs are foundational elements in the realms of the semantic web and knowledge management, providing structured ways to represent knowledge about the world. Defined as “a formal, explicit specification of a shared conceptualisation” by Noy and Hafner (1997), ontology, is a formal representation of knowledge as a set of concepts within a domain, and the relationships between those concepts (Gruber, 1993). It also provides standard vocabularies for researchers to annotate, share, and reuse their knowledge across different disciplines (Chen et al., 2018). Knowledge graphs further extend this concept by creating large networks of entities and their interrelations, encapsulated in graph structures, which enable complex queries and analytics, making knowledge more accessible and actionable (Ji et al., 2021).
2.1 Knowledge Representation for Smart Cities
@@
In smart city applications, ontology and knowledge graphs play pivotal roles in enhancing decision support and optimizing urban system. They facilitate the integration of diverse data sources, provide a basis for interoperability between different systems and platforms, and support the reasoning processes needed for intelligent decision-making (Chen et al., 2018). By mapping the complex relationships between various urban elements, such as transportation networks, utilities, and services, these knowledge representations synthesize data into actionable insights, thereby improving efficiency, sustainability, and the quality of urban life. (Bibri and Krogstie, 2017).
@@
From the technical perspective, ontology and knowledge graphs are crucial in the development of web-based decision support systems (Spoladore and Pessot, 2021), digital twins (Petrova-Antonova and Ilieva, 2021), databases (Balasubramani et al., 2016), AI applications (Bergman et al., 2018), and software dedicated to promote smart city management (Shi et al., 2023). Ontologies, with their hierarchical knowledge representations, provide the underlying architecture to structure and connect these technologies into meaningful software components and workflows, dynamically modeling and simulating the real-world urban environment (Al-Bayati, 2019). This configuration enables cities to anticipate problems, optimize operations, and engage more effectively with citizens (Kuster et al., 2020). Knowledge graphs, derived from ontologies, are topological networks that represent real-world entities (e.g., objects, events, situations, or concepts) and illustrate the relationships between them (Simsek et al., 2022). These networks are more application-focused in the context of data integration, management, and retrieval systems. They are widely applied in urban data applications to facilitate semantic search, recommendation systems, and AI models to link and contextualize disparate data (Chaudhri et al., 2022; Syed et al., 2022). Through the use of these semantic technologies, smart city applications, such as digital twin and online decision support systems, can evolve into more adaptive, responsive, and intelligent software applications, harnessing the full potential of the online big urban data to meet the challenges of complex urban system management and optimization (Weil et al., 2023b).
2.2 Ontology Construction
Methodologies for constructing scientific ontologies often incorporate both detailed and systematic approaches to effectively address the multifaceted nature of ontology development (Cristani and Cuel, 2005). At the micro level, these methodologies are crucial for focusing on the intricate details of ontology creation, such as the precise formalization of axioms and the design decisions shaped by the selected representation language (Guizzardi, 2005). This attention to detail ensures the accuracy and consistency of the ontology, which are vital for its reliability and usability. Conversely, macro-level approaches are essential for providing a holistic systems perspective, allowing for a comprehensive understanding of the development process through the lens of information systems and IT (Merali, 2006). By considering the broader context, these approaches facilitate the integration and alignment of the ontology with existing systems and processes, ensuring its scalability and adaptability. Together, these methodologies offer a balanced and thorough approach to ontology development, accommodating both specific technical considerations and overarching structural coherence. Well-known methodologies for constructing ontologies include older frameworks like Methontology and On-To-Knowledge, alongside more contemporary ones that focus on the re-usability and flexibility of ontology, such as NeON.
@@ 2.2 Ontology Construction
    The On-To-Knowledge methodology (OTKM) represents a structured approach to ontology design, aimed at enhancing knowledge management within organizations. It focuses on converting implicit company knowledge into an explicit, structured form, facilitating better decision-making and process optimization (Sure et al., 2004). This method emphasizes early stakeholder involvement, iterative development, and the use of formal ontologies to ensure that the captured knowledge is both accurate and usable in automated systems.
@@
    The NeON methodology caters to the evolving demands of ontology engineering, supporting the reuse, re-engineering, and integration of ontological resources (Gómez-Pérez and Suárez-Figueroa, 2009). By offering a diverse set of strategies and tools, NeON facilitates the development of robust ontologies and ontology networks, addressing complex semantic frameworks (Suárez-Figueroa et al., 2015). This methodology is particularly beneficial for projects requiring collaborative and distributed development, making it a pivotal resource for advancing semantic technologies and their applications in various domains.
@@
2.3 Traditional Methods for Ontology Implementation
At the implementation level, a variety of software tools for ontology creation, visualization, manipulation, and editing are widely utilized by knowledge engineers and domain scientists alike. Among these, Protégé (Noy et al., 2001), OntoSoft (Gil et al., 2016), Hozo (Mizoguchi et al., 2007), and OntoGraf (Falconer, 2010), yEd Graph Editor (Sedlmeier and Gogolla, 2016), Visual Understanding Environment (VUE) (Saigal and Kumar, 2005), and CmapTools (Cañas et al., 2004) have emerged as notable options in recent years. When choosing these tools, ontology developers must consider several critical factors, including the capability for reusing existing ontologies, comprehensive documentation, as well as the ability to export and import data across diverse formats, views, and libraries (Dudáš et al., 2018). These considerations are essential for ensuring the effective application and integration of ontological tools in scientific and engineering workflows.
Implementing ontology with specialized management software is a time-consuming and complex process, largely due to the detailed work required to represent domain-specific knowledge accurately and the expertise needed to effectively utilize these tools (Botzenhardt et al., 2011). The development of ontology is inherently iterative, involving repeated cycles of testing, evaluation, and refinement. Each cycle can reveal new requirements or inconsistencies, necessitating a return to previous stages, which extends the duration of the project (Fernández-López and Gómez-Pérez, 2002). Given these challenges, there is a pressing need to harness emerging AI technologies to automate the ontology construction, potentially streamlining the process and reducing the manual effort involved.
@@ 2.4 Semantic Web Standards
After being implemented, scientific knowledge representations are typically stored and documented using common semantic web languages that adhere to the semantic web standards, including RDF (Resource Description Framework), OWL (Web Ontology Language), and SPARQL (SPARQL Protocol and RDF Query Language) (Allemang and Hendler, 2011). These standards are crucial for structuring, storing, and querying data effectively. They enable the structured representation and interconnected querying of data across diverse systems, facilitating seamless data integration and accessibility on the Semantic Web.
@@
Together, these semantic web languages form the backbone of semantic web applications, allowing for robust and sophisticated handling of scientific resources, such as the sharing of scientific datasets and models, across the internet. Many smart city applications depend on semantic web technologies to discover and integrate essential data for promoting a variety of smart city services. Consequently, these technologies play a pivotal role in the development of numerous urban decision support systems and digital twins. Therefore, it is imperative that future AI-powered techniques for constructing ontologies adhere to established Semantic Web Standards. This alignment will ensure compatibility and enhance the effectiveness of data integration across diverse platforms and systems.
@@ 2.5 AI-powered Autonomous System
The swift evolution of AI technologies has catalyzed the emergence of autonomous systems(Brustoloni, 1991), which are sophisticated computer systems designed to undertake tasks and make decisions autonomously, with little to no human oversight (Wooldridge and Jennings, 1995). These agents have demonstrated significant potential across diverse fields such as robotics, healthcare, transportation, and finance, showcasing benefits like enhanced efficiency, diminished errors, and the capability to manage intricate tasks that might be impractical or exceedingly laborious for humans (Stone et al., 2022). Diverging from automated systems, such as automated irrigation systems and home automation systems, which execute predefined tasks based on specific sets of commands, inputs, rules, and environmental conditions, autonomous systems exhibit higher complexity, better adaptability to their surroundings, and the ability to make informed decisions and content generation from collected data (Albrecht and Stone, 2018).
@@
The capability of most existing rule- or algorithm-based reasoning cores is confined to certain or closed environments, limiting their reasoning capacity for tasks that involve creating new content or computer codes. With the advancement of generative AI models, LLMs, such as GPT-3 (Brown et al., 2020) and GPT-4 (Sun et al., 2023), has significantly propelled forward the research and development of autonomous systems. These models exhibit a profound comprehension of human natural language, enabling them to adeptly perform tasks across various domains including reasoning, creative writing, code generation, translation, and information retrieval (Li and Ning, 2023). With recent advancements in prompt engineering and tuning techniques, LLMs are increasingly being integrated as the reasoning core to enable AI-powered assistants capable of performing specialized tasks, such as generating content within specific domain contexts, automating code generation, and synthesizing data (Shin et al., 2023). This integration highlights the transformative impact of LLMs in enhancing automated processes and driving innovation across various fields, shedding light on an autonomous approach for generating scientific ontologies.
2.6 LLM-powered Knowledge Engineering
In recent years, there is an emerging trend to explore the integration of Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG) techniques with knowledge representations. This integration aims to enhance the accuracy and reliability of generative AI models, particularly when they are tasked with answering specialized scientific questions (Meyer et al., 2023; Baldazzi et al., 2023). As a majority of current studies focus on utilizing existing ontologies and knowledge graphs to improve the response accuracy of large language models (LLMs) and reduce potential hallucinations, there are only a few ongoing studies, most of which are presented as arXiv preprints, that have explored the potential of LLMs for constructing ontologies and knowledge graphs.
Babaei Giglou et al. (2023) introduces LLMs4OL, leveraging LLMs for Ontology Learning (OL). The study evaluates nine LLM models using zero-shot prompting for term typing, taxonomy discovery, and non-taxonomic relation extraction across domains like WordNet, GeoNames, and UMLS. Results show that while foundational LLMs alone may not suffice for complex ontology construction, fine-tuning can alleviate knowledge acquisition bottlenecks. Zhang and Soh (2024) presents the Extract-Define-Canonicalize (EDC) framework for automating knowledge graph creation using LLMs. EDC overcomes schema context limitations by splitting the process into open information extraction, schema definition, and canonicalization, handling both predefined and auto-constructed schemas. It outperforms state-of-the-art methods in extracting high-quality triplets across benchmark datasets. Kommineni et al. (2024) explores LLMs in automating ontology and knowledge graph construction using , introducing a semi-automated pipeline that minimizes human involvement while maintaining quality. This method, applied to deep learning methodologies, shows that LLMs can effectively assist with reduced human effort, though human validation remains crucial. Caufield et al. (2024) describes SPIRES, a method using LLMs for zero-shot learning and recursive prompt interrogation to automate knowledge base and ontology construction. SPIRES ensures accurate grounding with existing ontologies and excels across various domains, offering flexibility and ease of customization. It supports combining human expertise with advanced AI techniques to convert unstructured text into structured knowledge and is available in the open-source OntoGPT package.
3 Knowledge Gaps and Contributions